Fred is 5 years older than Mary. In 4 years, Fred will be twice as old as Mary was one year ago. What is Fred's age now?
Answer by Alan3354(69443)

Fred is 5 years older than Mary. In 4 years, Fred will be twice as old as Mary was one year ago. What is Fred's age now?
-----------------
F = M + 5
--------
In 4 years, Fred will be twice as old as Mary was one year ago.
---
F + 4 = 2*(M-1)
Can you do the rest?